OpenStudy (anonymous):

system of equations, solve for x and y, 3x^2-2y and -3y^2-2x

OpenStudy (dumbcow):

where are the = signs

OpenStudy (anonymous):

3x^2-2y=0 and -3y^2-2x=0

OpenStudy (llort):

(0, 0) (-2/3, 2/3)

OpenStudy (anonymous):

LLort I found (0,0) but how did you get (-2/3,2/3)?

OpenStudy (llort):

I graphed it

OpenStudy (dumbcow):

solve first one for y y =(3x^2)/2 substitute it into 2nd equ -3(9x^4/4) -2x = 0 multiply everything by 4 -27x^4 -8x = 0 -x(27x^3 +8) = 0 -x(3x+2)(9x^2 -6x +4) = 0 real solutions x=0, x = -2/3 substitute in to find y y =0, y = 2/3

OpenStudy (anonymous):

dumbcow let me take a look at this, be a minute.

OpenStudy (dumbcow):

ok there are 4 solutions (0,0) (0,2/3) (-2/3,0) (-2/3,2/3)

OpenStudy (anonymous):

How did 27x^3+8) become (3x+2)(9x^2 -6x +4)?

OpenStudy (dumbcow):

oh sorry, thats the sum of cubes formula (a+b)^3 = (a+b)(a^2 -ab+b^2) 27x^3 = (3x)^3 8 = 2^3

OpenStudy (dumbcow):

*(a^3 +b^3)

OpenStudy (anonymous):

I remember now. Thanks, the 2/3 is what i was looking for.